OpenStudy (anonymous):

Find the exact circumference of a circle with an area equal to 36pi sq. in. a) 12pi b) 18pi c) 324pi

OpenStudy (anonymous):

@satellite73

TheSmartOne (thesmartone):

what is the formula for area and circumference?

OpenStudy (anonymous):

circumference of a circle is \(C=2\pi r\) where \(r\) is the radius

OpenStudy (anonymous):

yeah that's what I was gonna' say

OpenStudy (anonymous):

area is \(\pi r^2=36\pi\) so \(r=?\)

OpenStudy (anonymous):

18

TheSmartOne (thesmartone):

wrong

OpenStudy (anonymous):

It is?

OpenStudy (anonymous):

not half dear, the square root of 36

OpenStudy (anonymous):

Oh!

OpenStudy (anonymous):

I'm sorry

OpenStudy (anonymous):

It's 6 then

OpenStudy (anonymous):

not \(36\div 2\) but rather \(\sqrt{36}\)

TheSmartOne (thesmartone):

area is \(\sf\pi r^2=36\pi \to r^2 = 26\)

OpenStudy (anonymous):

right, and if \(r=6\) then your final job is to compute \[2\times 6\times \pi\]

TheSmartOne (thesmartone):

and now plug in \(\sf r=6\) into the circumference formula (:

OpenStudy (anonymous):

37?

OpenStudy (anonymous):

For 2*6*3014=37

OpenStudy (anonymous):

lol no

TheSmartOne (thesmartone):

You don't simplify the \(\sf\pi\) since it wants the exact circumference

OpenStudy (anonymous):

just \[2\times 6\pi\]what is \(2\times 6\)?

OpenStudy (anonymous):

12

TheSmartOne (thesmartone):

and the answer choices also contain \(\sf \pi\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

right, \(12\pi\)

OpenStudy (anonymous):

Omg it's 12pi

OpenStudy (anonymous):

omg indeed!

TheSmartOne (thesmartone):

\(\ddot\smile\)

OpenStudy (anonymous):

I'm sorry after spending a week traveling up and down from Mississippi to Florida, My head hasn't been on straight

OpenStudy (anonymous):

that is what you get for going to mississippi

OpenStudy (anonymous):

Yeah, but I was helping my aunt on bisiness and plus I made some money

OpenStudy (anonymous):

ok ok no offence

OpenStudy (anonymous):

*business* Lol

OpenStudy (anonymous):

$$ always good

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

bidniz as we say

OpenStudy (anonymous):

btw sent you lee morgan youtube again take a listen some time

TheSmartOne (thesmartone):

x'D

OpenStudy (anonymous):

I haven't made any in awhile and plus it's always better thaqn spending a week alone with my gma and uncle bickering a complaining the entire time

OpenStudy (anonymous):

Don't worry I will. It's just that I'm at school right now and their wifi blocks youtube

OpenStudy (anonymous):

real school not on line? in the summer? sounds like a drag

OpenStudy (anonymous):

Yeah. I have Odyssey

OpenStudy (anonymous):

I go home at 11pm

OpenStudy (anonymous):

OMG indeed !

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

But that's what I get after failing the last half of geometry. although most of it wasn't my fault after the difficult home, living, and family issues I was having....

OpenStudy (anonymous):

It's hard when you've been to three different highschools during your freshmen year

OpenStudy (anonymous):

you'll get through geometry

OpenStudy (anonymous):

i bet

OpenStudy (anonymous):

Thanks. I sure hope so. I'll still be taking Algebra 2 nxt year. I just have to complete Geometry online during the summer though

OpenStudy (anonymous):

lets get busy !

OpenStudy (anonymous):

got another ?

OpenStudy (anonymous):

Yep. I sure do

OpenStudy (anonymous):

Hey, which one of you should I medal?

OpenStudy (anonymous):

You both were alot of help

OpenStudy (anonymous):

smart one i got plenty

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

I will

OpenStudy (anonymous):

Thanks for making the decision easy

TheSmartOne (thesmartone):

I have plenty as will, give it to satellite so he can keep his 100 SS (:

OpenStudy (anonymous):

Oh mouse dung!

OpenStudy (anonymous):

tOO LATE TO GO BACK NOW

OpenStudy (anonymous):

Woops caps locked again

OpenStudy (anonymous):

I hate when that happens

OpenStudy (anonymous):

fret not i can't turn them in to cash or valuable prizes anyways

OpenStudy (anonymous):

Oh cool

OpenStudy (anonymous):

Well lets get on to the next question then

OpenStudy (anonymous):

good idea